OpenStudy (anonymous):

A store received $823 from the sale of 5 tape recorders and 7 radios. If the receipts from the tape recorders exceeded the receipts from the radios by $137, what was the cost of a tape recorder?

OpenStudy (anonymous):

let the cost of tape be x and that of radio be y. so 5x+7y=823............(1); 5x=137+7y...........(2) from 1 & 2 14y=686; or y=$49; hence x=(137+7*49)/5=96. please check it

OpenStudy (anonymous):

Thank you. I made a mistake with the elimination process. I was not getting the answer on the key. Now I see what I did wrong. Thank you.

OpenStudy (anonymous):

you welcome.